What is the pH change of a 0.240 M solution of citric acid (pKa=4.77) if citrate is added to a concentration of 0.160 M with no change in volume?
Answer #1

Since -

        pH = pKa + log [Citrate / citric acid]

             = 4.77 + log (0.160/0.240)

             = 4.59

Initail pH = -log (Ka*[H+])1/2 = (10-4.77*0.240)0.5 = -log (0.0020) = 2.7

pH change = 4.6-2.7 = 1.9